The combined area of a square and a rectangle is 116 square centimeters. The width of the rectangle is 2 centimeters more than the length of a side of the square, and the length of the rectangle is 2 centimeters more than its width.Find the dimensions of the square and the rectangle?
Answer by ankor@dixie-net.com(22740)

The combined area of a square and a rectangle is 116 square centimeters.
The width of the rectangle is 2 centimeters more than the length of a side of the square, and the length of the rectangle is 2 centimeters more than its width.
Find the dimensions of the square and the rectangle?
:
x = the side of the square and the width of the rectangle
:
Adding the two areas
x^2 + x(x+2) = 116
x^2 + x^2 + 2x = 116
A quadratic equation
2x^2 + 2x - 116 = 0
Solve this using the quadratic formula: a=2; b=2; c=-116